A taxi driver claims that his average monthly income is Php 2, 800.00 with a standard deviation of Php 250.00 . A sample of 25 drivers were surveyed and found that their average monthly income is Php 3,500.00with a standard deviation of Php 350.00.Test the hypothesis at 1% level of significance.

1
Expert's answer
2022-05-30T07:41:11-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=2800"

"H_1:\mu\not=2800"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a two-tailed test is "z_c = 2.5758."

The rejection region for this two-tailed test is "R = \{z:|z|>2.5758\}."

The z-statistic is computed as follows:



"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{3500-2800}{250/\sqrt{25}}=14"

Since it is observed that "|z|=14>2.5758=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(z>14)= 0," and since "p= 0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 2800, at the "\alpha = 0.01" significance level.